OpenStudy (anonymous):

8.9 + 1.2 (3a - 1) = 14.9

OpenStudy (anonymous):

1.2(3a-1) = 14.9-8.9 = 6 3a-1 = 6/1.2 =5 3a = 5+1 = 6 a = 6/3 = 2